OpenStudy (anonymous):

A 15 gram sample of radioactive iodine decays in such a way that the mass remaining after t days is given by m(t)=15e^(-0.087t), where m(t) is measures in grams. After how many days is there only 8 grams remaining?

OpenStudy (amistre64):

8 --- = e^(-.087t) 15 8 ln(---) = -.087t 15 ln(8/15) ------- = t -.087

OpenStudy (amistre64):

t = 7.225....

OpenStudy (amistre64):

dbl chk and we get: m(t) = 7.999999....

OpenStudy (anonymous):

Yayyyy! Thanks again :)

OpenStudy (amistre64):

youre welcome :)